# Accelerating Ray Tracing-Based Wireless Channels Generation for Real-Time Network Digital Twins

### v1

@@ Abstract
RT simulation is a widely used approach to enable modeling wireless channels in applications such as network digital twins. However, the computational cost to execute ray tracing (RT) is proportional to factors such as the level of detail used in the adopted 3D scenario. This work proposes RT pre-processing algorithms that aim at simplifying the 3D scene without distorting the channel. It also proposes a post-processing method that augments a set of RT results to achieve an improved time resolution. These methods enable using RT in applications that use a detailed and photorealistic 3D scenario while generating consistent wireless channels over time. Our simulation results with different 3D scenarios demonstrate that it is possible to reduce the simulation time by more than 50% without compromising the accuracy of the RT parameters.
@@ I Introduction
The remainder of this article is organized as follows. Section II describes the fundamental properties of the problems addressed, considering all the definitions and mathematical elements regarding a RT simulation and the channel model adopted. Section III positions the two problems in the context of network digital twins, illustrating how their solutions can be integrated. With the required nomenclature and problems defined, Section IV discusses the previously published related work, which helps localizing the proposed solutions with respect to the state-of-the-art. Section V presents the solutions to each of the posed problems, detailing the preprocessing and post-processing methods. Section VI separately validates the two proposed solutions in different scenarios with respect to channel characteristics and simulation time. In Section VII we conclude this article and point out future directions.
@@ II Problems Formulation
This article focuses on two distinct problems related to speeding up the generation of a sequence of wireless channels represented by MPC (or rays), as the ones obtained by conventional RT algorithms. The first problem consists of using post-processing to obtain an upsampled version of the RT results with improved time resolution. The goal is to achieve these results in significantly less time than running a conventional ray tracer software with improved time resolution, while still obtaining accurate wireless channels.
The second problem involves implementing a preprocessing stage to simplify 3D scenarios by reducing the polygon count while preserving the integrity of the respective sequence of wireless channels. For concreteness, it is assumed here that each 3D object is represented by a single mesh. [^1] For instance, organized a file of type .ply (polygon file format). Other representations can be used, but in the article paper, a 3D scene with $O$ objects is represented by a set $\cal S$ with $O$ meshes. The primary goal of these solutions is to accelerate RT simulations. Besides, such preprocessing methods may be crucial for enabling RT in cases where the complexity of the 3D scenario exceeds the capabilities of the RT software, and a simpler 3D scenario becomes mandatory.
@@
where a ray tracer calculates a set of rays $\mathcal{R}$ representing the wireless channel, given the 3D scene described by meshes in $\mathcal{S}$. It is then assumed that $\mathcal{R}$ provides all information needed to obtain the desired representation of the wireless channel. For instance, the experiments in this article assume that $\mathbf{H}$ is a narrowband multiple-input multiple-output (MIMO) channel matrix [21], obtained via the geometric model applied to the results of a single-input single-output (SISO) RT simulation [22]. But alternative signal processing could transform the rays into a wideband channel for simulating orthogonal frequency division multiplexing (OFDM) systems or a MIMO-OFDM channel.
It should be noted that the two posed problems can be considered supervised, in the sense that it is assumed there is a “ground-truth” to be used in quantitative (not only qualitative) comparisons. This ground-truth is obtained by invoking the conventional ray tracer for each scene ${\cal S}_{m}^{g}$ in the sequence $\widetilde{\mathbf{S}}_{\textrm{g}}=[{\cal S}_{m}^{g}]_{m=1}^{M}$ with the highest time resolution (that corresponds to a sampling interval $T_{\textrm{s}}$). This ground truth must be properly interpreted, especially because distinct ray tracers may lead to very different results [14]. It is out of the scope of this article to assess how accurate the RT results are with respect to measurements [23]. The following subsections describe the two problems.
@@ II-A Post-processing RT for improved time resolution
Figure 1: An illustrative description of the adopted nomenclature. The “blocks” represent brick walls that impose interactions across the ray trajectories. In this case, we have at discrete time $n=1$, a set of rays $\mathcal{R}_{1}$ with $K_{1}=K_{\textrm{{max}}}=3$. The ray $R_{3}$ represents a LOS ray; $R_{1}$ and $R_{2}$ are NLOS. For each ray $R$, the 3-tuple with p, $b$, and $\mathcal{I}$ is calculated by the ray tracer.
@@
Each ray $R=(\mathbf{p},b,\mathcal{I})$ is a tuple of three elements: a vector $\mathbf{p}\in\mathbb{R}^{7}$ with the numerical parameters listed in Table I; a binary variable $b$ indicating whether the ray is LOS or NLOS, and a set $\mathcal{I}$ describing all interactions between the ray and the objects composing the 3D scene. And each element of the set $\mathcal{I}$ is a tuple with two elements: a vector $\mathbf{x}=(x,y,z)$ informing the position in meters of the interaction, with respect to the scene origin $(0,0,0)$; and an integer $t$ indicating the type of propagation phenomenon (reflection, diffraction, diffusion scattering, refraction / transmission, etc.) undergone by the ray at the given interaction.
@@ II-B Preprocessing RT 3D input
Another problem investigated in this article is how to enable or speed up RT by simplifying the 3D scenes without compromising accuracy and important features of channel sequences such as spatial consistency [26]. Engines such as Unreal [27] and Unity [28] support scenes with a number of faces that are orders of magnitude larger than the number supported by existing ray tracer software. Hence, in spite of solutions to P1 using RT augmentation, executing RT itself can be unfeasible for extremely detailed 3D scenarios.
@@ II-C Multi-objective framework to define the two problems
Section V describes solutions to problems P1 and P2. Before that, the next section discusses in a high level how both solutions can be used in network digital twins.
III Speeding Up RT in Network Digital Twins
@@
The methods in the category cut-out-based simplification are novel and described in Section V. The main idea is that, as the simulation evolves over time, the TRP and scatterers eventually move to new positions and the SOH module can reduce the effective volume (or area) to be simulated based on previous RT runs and new object positions. The SOH can then create and send to the ray tracer a simpler 3D scene.
IV Related Work
@@
The work closest to ours is the one by Zentner and Mucalo [13], which formulated an application that uses an interpolation method in the ray parameter domain, leveraging the idea of ray entities that groups a set of similar rays that undergo the same propagation phenomenon across the interactions with the objects in the scenario. Organizing similar rays helped interpolation methods such as cubic splines, linear, or polynomial, to augment the sequence originally generated by the conventional ray tracer as suggested in problem P1. Our solution to P1 has the following improvements with respect to [13]: a) we consider all parameters that compose a set of rays $\cal R$ while the authors in [13] did not consider the ray phase (which is the hardest parameter to estimate), but only the time of arrival, angles of departure and elevation, and received power, and b) we consider the existence of a ray birth-and-death process that was not taken into account in [13].
@@
The mentioned solutions to P2 are based on mesh simplification. Alternative approaches were adopted by Mi et al. [33], which used two methods called uniform grid space partitioning and $k$-D tree space. Both methods share some similarities to the cut-out based methods proposed in this article. The former method divides the environment into grids and discards the ones (and the objects that belong to this grid) that do not interact with the RT paths, reducing the number of “intersection tests”. The latter method also uses the idea of grid division, but in this case considers a non-uniform partitioning based on $k$-D tree algorithm [31]. A main drawback related to these methods is the need to divide the 3D scenario to check which grid is not contributing to the RT simulation. In large scenarios, this process can consume a considerable amount of time.
@@ V Proposed Solutions
This section describes our solutions to the problems P1 and P2. The method Augmented Ray-Tracing poSt-processing (ARTS) can be implemented in the WCG module of a digital twin to solve P1. After the discussion of ARTS, methods to solve P2 are then presented.
@@ V-A Augmented Ray-Tracing Post-processing (ARTS) method
The input-output relationships in our method are described in Fig. 5. The core of this implementation is represented by the ARTS block and is based on three main steps, which are based on the concepts of pairing rays, equivalent rays, and stable segments. This comparison for pairing and checking ray equivalence relies on a corresponding face identification (ID). The faces of all objects composing the scene are uniquely identified, and the interaction position $(x,y,z)$ allows one to obtain the face ID even in case the ray tracer does not directly provide it. Hence, the extended interactions set $I_{r}^{n}$ of ray $R_{r}^{n}$ is denoted as $\mathcal{I}_{r}^{n}=\{(\mathbf{x}_{r}^{n}[1],t^{n}_{r}[1],d^{n}_{r}[1]),\ldots,(\mathbf{x}_{r}^{n}[I_{r}^{n}],t^{n}_{r}[I_{r}^{n}],d^{n}_{r}[I_{r}^{n}])\}$, where $d$ is an integer number.
@@
To achieve successful ray generation with small error, the ray augmenter (ARTS block) leverages spatial consistency between scenes. In other words, before generating unknown ray parameters, each known ray—equivalent to another—is paired to form a stable segment. This pairing is essential because the ray tracer does not preserve the index of each ray throughout the simulation over time. The detailed structure of the ARTS method is outlined in Algorithm 1.
@@
4:$\tilde{\mathbf{R}}_{\text{a}}\leftarrow[\mathcal{R}^{\prime}_{1},\mathcal{R}^{\prime}_{2},\dots,\mathcal{R}^{\prime}_{M}]$triangleright Initialize empty upsampled sequence
@@
16:for $n=1$ to $N-1$ $\triangleright$ Iterate over scenes
@@
21:                  $P^{p}_{r}\leftarrow\frac{p-1}{U}(R^{n+1}_{r}-R^{n}_{r})$
@@
25:                       $P^{p}_{r}\leftarrow R^{n}_{r}$triangleright Deals with dead ray
@@
27:                       $P^{p}_{r}\leftarrow R^{n+1}_{r}$triangleright Deals with born ray
@@
30:              $\mathcal{R}^{\prime}_{p}\leftarrow\mathcal{R}^{\prime}_{p}\cup\{P_{r}^{p}\}$triangleright Update set with a new ray
@@
An example of it can be visualized in Fig. 6, hypothetically considering that scene 2 is unknown, and we would like to generate the LOS ray (in green) in this scene. We will find that the LOS ray in scenes 1 and 3 is composed of stable segments since all requirements are accomplished. With this equivalence verified, we can use a linear interpolation method (line 21) defined by
$P^{p}_{r}=\frac{p-1}{U}(R^{n+1}_{r}-R^{n}_{r}).$ (3)
When the equivalence is not detected between two rays (line 22) to form a stable segment (for instance, the NLOS red ray in Fig. 6), this is depicted as a case of dead or born ray [41], which means that there exists an untraceable ray in the unknown scene that was born with a new characteristic or died in this exact scene, as depicted in Fig. 6. Therefore, as it is not possible to know exactly if a ray will die or be born, we can model this behavior with different levels of complexity [41]. In the ARTS method, we adopt a simplified approach, considering that an untraceable ray can be modeled via a random variable $B$ obeying a Bernoulli distribution (line 23), $B\sim\mathcal{B}(\ell)$, with $\ell=0.5$. So, given that a ray is untraceable, we have an equal probability of having a born or dead ray in the scene to be generated.
Therefore, given an untraceable ray $P^{p}_{r}$ is detected in the $p$-th scene, if it was assigned a dead ray, this ray will assume characteristics from the $r$-th ray from the last known scene (line 25). Otherwise, if it was assigned as a born ray, this ray will assume the characteristics from the $r$-th ray from the next scene (line 27).
@@ V-B1 Mesh-based methods
In our digital twin framework, two mesh-based algorithms are implemented: vertex clustering and quadric edge collapse [43].
The vertex clustering algorithm operates on a grid of pixels (for 2D meshes) or voxels (for 3D meshes) of size $V$. To illustrate this, consider a $3\times 3$ grid applied to a 2D mesh, as shown in Fig. 8(1). The same concept extends to 3D meshes without loss of generality. Given a cell size determined by $\varepsilon$, which corresponds to an area of $\varepsilon^{2}$ (11.1% of the original pixel size), the algorithm aims to reduce the vertex count by clustering multiple vertices within each grid cell and assigning a representative vertex $L$.
In other words, if a cell contains multiple vertices, they are mapped to a single representative vertex. This vertex is computed either as the weighted average of the coordinates of all vertices in the cell or as the position of the most heavily weighted vertex. The weights are determined based on two criteria [44, 45]: (i) the probability that a vertex lies on the object’s silhouette from a given viewpoint, and (ii) the presence of a large face associated with the vertex. A vertex that satisfies both criteria—having a high silhouette probability and being part of a large face—is assigned a higher weight.
@@
$L=\frac{1}{Q}\sum\limits_{i=1}^{Q}(Wv_{x_{i}},Wv_{y_{i}},Wv_{z_{i}}),$ (4)
where $W$ is the weight of the vertex. Hence, given the evaluated representative vertex, the simplified mesh looks like Fig. 8(2).
@@
Figure 9: Illustration of the quadric edge collapse algorithm applied to a 2D mesh. In (1), the dotted edge represents the edge selected for collapse, while the hashed triangles indicate the affected regions. The two arrows pointing toward each other signify the vertices that will be merged. In (2), the resulting 2D mesh is shown after the edge collapse operation.

### v2

@@ Abstract
RT simulation is a widely used approach to enable modeling wireless channels in applications such as network digital twins. However, the computational cost to execute ray tracing (RT) is proportional to factors such as the level of detail used in the adopted 3D scenario. This work proposes RT pre-processing algorithms that aim at simplifying the 3D scene without distorting the channel, by reducing the scenario area and/or simplifying object shapes in the scenario. It also proposes a post-processing method that augments a set of RT results to achieve an improved time resolution. These methods enable using RT in applications that use a detailed and photorealistic 3D scenario while generating consistent wireless channels over time. Our simulation results with different urban scenarios scales, in terms of area and object details, demonstrate that it is possible to reduce the simulation time by more than 50% without compromising the accuracy of the multipath RT parameters, such as angles of arrival and departure, delay, phase, and path gain.
@@ I Introduction
The remainder of this article is organized as follows. Section II describes the fundamental properties of the problems addressed, considering all the definitions and mathematical elements regarding a RT simulation and the channel model adopted. With the required nomenclature and problems defined, Section III discusses the previously published related work, which helps localizing the proposed solutions with respect to the state-of-the-art. Section IV positions the two problems in the context of network digital twins, illustrating how their solutions can be integrated. Section V presents the solutions to each of the posed problems, detailing the preprocessing and post-processing methods. Section VI separately validates the two proposed solutions in different scenarios with respect to channel characteristics and simulation time. In Section VII we conclude this article and point out future directions.
@@ II Problems Formulation
This article focuses on two distinct problems related to speeding up the generation of a sequence of wireless channels represented by MPCs (or rays), as the ones obtained by conventional RT algorithms. The first problem consists of using post-processing to obtain an upsampled version of the RT results with improved time resolution. The goal is to achieve these results in significantly less time than running a conventional ray tracer software with improved time resolution, while still obtaining accurate wireless channels.
The second problem involves implementing a preprocessing stage to simplify 3D scenarios by reducing the polygon count while preserving the integrity of the respective sequence of wireless channels. For concreteness, it is assumed here that each 3D object is represented by a single mesh. [^1] For instance, organized a file of type .ply (polygon file format). Other representations can be used, but in this article, a 3D scene with $O$ objects is represented by a set $\cal S$ with $O$ meshes. The primary goal of these solutions is to accelerate RT simulations. Besides, such preprocessing methods may be crucial for enabling RT in cases where the complexity of the 3D scenario exceeds the capabilities of the RT software, and a simpler 3D scenario becomes mandatory.
@@
where a ray tracer calculates a set of rays $\mathcal{R}$ representing the wireless channel, given the 3D scene described by meshes in $\mathcal{S}$. It is then assumed that $\mathcal{R}$ provides all information needed to obtain the desired representation of the wireless channel. For instance, the experiments in this article assume that $\mathbf{H}$ could assume a narrowband (NB) or wideband (WB) multiple-input multiple-output (MIMO) channel matrix [21], obtained via the geometric model applied to the results of a single-input single-output (SISO) RT simulation [22].
It should be noted that the two posed problems can be considered supervised, in the sense that it is assumed there is a “ground-truth” to be used in quantitative (not only qualitative) comparisons. This ground-truth is obtained by invoking the conventional ray tracer for each scene ${\cal S}_{m}^{\textrm{g}}$ in the sequence $\widetilde{\mathbf{S}}_{\textrm{g}}=[{\cal S}_{m}^{\textrm{g}}]_{m=1}^{M}$ with the highest time resolution (that corresponds to a sampling interval $T_{\textrm{s}}$). This ground truth must be properly interpreted, especially because distinct ray tracers may lead to very different results [14]. It is out of the scope of this article to assess how accurate the RT results are with respect to measurements [23]. The following subsections describe the two problems.
@@ II-A Post-processing RT for improved time resolution
Figure 1: An illustrative description of the adopted nomenclature. The “blocks” represent brick walls that impose interactions across the ray trajectories. In this case, we have at discrete time $n=1$, a set of rays $\mathcal{R}_{1}$ with $K_{1}=K_{\textrm{max}}=3$. The ray $R_{3}$ represents a LOS ray; $R_{1}$ and $R_{2}$ are NLOS. For each ray $R$, the 3-tuple with $\mathbf{p}$, $b$, and $\mathcal{I}$ is calculated by the ray tracer.
@@
Each ray $R=(\mathbf{p}^{n}_{r},b^{n}_{r},\mathcal{I}^{n}_{r})$ is a tuple of three elements: a vector $\mathbf{p}^{n}_{r}\in\mathbb{R}^{7}$ with the numerical parameters listed in Table I; a binary variable $b^{n}_{r}$ indicating whether the ray is LOS or NLOS, and a set $\mathcal{I}^{n}_{r}$ describing all interactions between the ray and the objects composing the 3D scene. And each element of the set $\mathcal{I}^{n}_{r}$ is a tuple with two elements: a vector $\mathbf{x}=(x,y,z)$ informing the position in meters of the interaction, with respect to the scene origin $(0,0,0)$; and an integer $t$ indicating the type of propagation phenomenon (reflection, diffraction, diffusion scattering, refraction / transmission, etc.) undergone by the ray at the given interaction.
@@ II-B Preprocessing RT 3D input
Another problem investigated in this article is how to enable or speed up RT by simplifying the 3D scenes without compromising accuracy and important features of channel sequences such as spatial consistency [26]. Engines such as Unreal [27] and Unity [28] support scenes with a number of faces that are orders of magnitude larger than the number supported by existing ray tracer software, as depicted in Fig. 2. Hence, in spite of solutions to P1 using RT augmentation, executing RT itself can be unfeasible for extremely detailed 3D scenarios.
Figure 2: Scenario called Modern City rendered with Unreal engine.
@@ II-C Multi-objective framework to define the two problems
Section V describes solutions to problems P1 and P2. Before that, the next section discusses the related works regarding the P1 and P2 problems.
III Related Work
@@
The work closest to ours is the one by Zentner and Mucalo [13], which formulated an application that uses an interpolation method in the ray parameter domain, leveraging the idea of ray entities that groups a set of similar rays that undergo the same propagation phenomenon across the interactions with the objects in the scenario. Organizing similar rays helped interpolation methods such as cubic splines, linear, or polynomial, to augment the sequence originally generated by the conventional ray tracer as suggested in problem P1. Our solution to P1 has the following improvements with respect to [13]: a) we consider all parameters that compose a set of rays $\cal R$ while the authors in [13] did not consider the ray phase (which is the hardest parameter to estimate), but only the time of arrival, angles of departure and elevation, and received power; and b) we consider the existence of a ray birth-and-death process that was not taken into account in [13].
@@
The mentioned solutions to P2 are based on mesh simplification. Alternative approaches were adopted by Mi et al. [31], which used two methods called uniform grid space partitioning and $k$-D tree space. Both methods share some similarities to the cut-out based methods proposed in this article. The former method divides the environment into grids and discards the ones (and the objects that belong to this grid) that do not interact with the RT paths, reducing the number of “intersection tests”. The latter method also uses the idea of grid division, but in this case, considers a non-uniform partitioning based on $k$-D tree algorithm [39]. A main drawback related to these methods is the need to divide the 3D scenario to check which grid is not contributing to the RT simulation. In large scenarios, this process can consume a considerable amount of time.
IV Speeding up RT in network digital twins
The methods in the category cut-out-based simplification are novel and described in Section V. The main idea is that, as the simulation evolves over time, the TRP and scatterers eventually move to new positions, and the SOH module can reduce the effective volume (or area) to be simulated based on previous RT runs and new object positions. The SOH can then create and send to the ray tracer a simpler 3D scene.
@@ V Proposed Solutions
This section describes our solution to the problems P1 and P2. The method Augmented Ray-Tracing poSt-processing (ARTS) can be implemented in the WCG module of a digital twin framework to solve P1. After the discussion of ARTS, methods to solve P2 are then presented.
@@ V-A Augmented Ray-tracing post-processing (ARTS) method
The input-output relationships in our method are described in Fig. 6. The core of this implementation is represented by the ARTS block and is based on three main steps, which are based on the concepts of pairing rays, equivalent rays, and stable segments.
@@
This comparison for pairing and checking ray equivalence relies on a corresponding face identification (ID). The faces of all objects composing the scene are uniquely identified, and the interaction position $(x,y,z)$ allows one to obtain the face ID even in case the ray tracer does not directly provide it. Hence, the extended interactions set $I_{r}^{n}$ of ray $R_{r}^{n}$ is denoted as $\mathcal{I}_{r}^{n}=\{(\mathbf{x}_{r}^{n}[1],t^{n}_{r}[1],d^{n}_{r}[1]),\ldots,(\mathbf{x}_{r}^{n}[I_{r}^{n}],t^{n}_{r}[I_{r}^{n}],d^{n}_{r}[I_{r}^{n}])\}$, where $d$ is an integer number. So, to achieve successful ray generation with small error, the ray augmenter (ARTS block) leverages spatial consistency between scenes. In other words, before generating unknown ray parameters, each known ray—equivalent to another—is paired to form a stable segment. This pairing is essential because the ray tracer does not preserve the index of each ray throughout the simulation over time. The detailed structure of the ARTS method is outlined in Algorithm 1.
@@
4:$\tilde{\mathbf{R}}_{\text{a}}\leftarrow[\mathcal{R}^{\prime}_{1},\mathcal{R}^{\prime}_{2},\dots,\mathcal{R}^{\prime}_{p},\dots,\mathcal{R}^{\prime}_{M}]$triangleright Initialize empty upsampled sequence
@@
16:for $n=1$ to $N-1$ $\triangleright$ Iterate over known scenes
@@
21:                  $\mathbf{p}^{p}_{r}\leftarrow\mathbf{p}_{r}^{n}+\frac{p-1}{U}(\mathbf{p}^{n+1}_{r}-\mathbf{p}^{n}_{r})$
@@
25:                       $\mathbf{p}^{p}_{r}\leftarrow\mathbf{p}^{n}_{r}$triangleright Deals with dead ray
@@
27:                       $\mathbf{p}^{p}_{r}\leftarrow\mathbf{p}^{n+1}_{r}$triangleright Deals with born ray
@@
30:         end for
@@
An example of it can be visualized in Fig. 7, hypothetically considering that scene 2 is unknown, and we would like to generate the LOS ray (in green) in this scene. We will find that the LOS ray in scenes 1 and 3 composes stable segments since all requirements are accomplished. With this equivalence verified, we can use a linear interpolation method (line 21) defined by
$\mathbf{p}^{p}_{r}=\mathbf{p}_{r}^{n}+\frac{p-1}{U}(\mathbf{p}^{n+1}_{r}-\mathbf{p}^{n}_{r}),$ (3)
where $1<p<U+1\hskip 4.26773pt\forall p\in\mathbb{Z}^{+}$.
When the equivalence is not detected between two rays (line 22) to form a stable segment (for instance the NLOS red ray in Fig. 7), this is depicted as a case of dead or born ray [41], which means that there exists an untraceable ray in the unknown scene that was born with a new characteristic or died in this exact scene, as depicted in Fig. 7. Therefore, as it is not possible to know exactly if a ray will die or be born, we can model this behavior with different levels of complexity [41]. In the ARTS method, we adopt a simplified approach, considering that an untraceable ray can be modeled via a random variable $B$ obeying a Bernoulli distribution (line 23), $B\sim\mathcal{B}(\ell)$, with $\ell=0.5$. So, given that a ray is untraceable, we have an equal probability of having a born or dead ray in the scene to be generated.
Therefore, given an untraceable ray $R^{p}_{r}$ is detected in the $p$-th scene, if it was assigned a dead ray, this ray will assume characteristics from the $r$-th ray from the last known scene (line 25). Otherwise, if it was assigned as a born ray, this ray will assume the characteristics from the $r$-th ray from the next scene (line 27).
@@ V-B1 Mesh-based methods
In our SOH module, two mesh-based algorithms are implemented: vertex clustering and quadric edge collapse [43]. The vertex clustering algorithm operates on a grid of pixels (for 2D meshes) or voxels (for 3D meshes) of size $V$. To illustrate this, consider a $3\times 3$ grid applied to a 2D mesh, as shown in Fig. 9(1). The same concept extends to 3D meshes without loss of generality. Given a cell size determined by $\varepsilon$, which corresponds to an area of $\varepsilon^{2}$ (11.1% of the original pixel size), the algorithm aims to reduce the vertex count by clustering multiple vertices within each grid cell and assigning a representative vertex $L$. In other words, if a cell contains multiple vertices, they are mapped to a single representative vertex. This vertex is computed either as the weighted average of the coordinates of all vertices in the cell or as the position of the most heavily weighted vertex. The weights are determined based on two criteria [44, 45]: (i) the probability that a vertex lies on the object’s silhouette from a given viewpoint, and (ii) the presence of a large face associated with the vertex. A vertex that satisfies both criteria—having a high silhouette probability and being part of a large face—is assigned a higher weight.
@@
$L=\frac{1}{Q}\sum\limits_{i=1}^{Q}(Yv_{x_{i}},Yv_{y_{i}}),$ (4)
where $Y$ is the weight of the vertex. Hence, given the evaluated representative vertex, the simplified mesh looks like Fig. 9(2).
@@
Figure 10: Illustration of the quadric edge collapse algorithm applied to a 2D mesh. In (1), the dashed edge represents the edge selected for collapse, while the hashed triangles indicate the affected regions. The two arrows pointing toward each other signify the vertices that will be merged. In (2), the resulting 2D mesh is shown after the edge collapse operation.